# Hierarchical Neural Simulation-Based Inference Over Event Ensembles

### v1

@@ 4 Conclusions
Appendix A Further details on experiments
@@ A.1 Details of ‘Simple multi-variate normal likelihood’ experiments
The estimators are trained using the AdamW [17, 15] optimizer with initial learning rate $3\times 10^{-4}$ and cosine annealing over 100 epochs. The checkpoint corresponding to the lowest validation loss is used. Evaluation is performed on 500 new test samples, evaluating the posterior estimator for sequences with cardinality varied from 1 to 200 for each test sequence.
@@ A.2 Details of ‘Mixture models in particle physics: frequentist treatment’ experiments
We use the implementation of the NUTS algorithm available in PyMC5 [25]. For each inference run, two Markov Chains of 50k samples each are sampled. The first 12k samples in each chain are discarded and not used for the estimation of the posterior.
@@ A.4 Details of ‘Astrophysics example: Strong gravitational lensing’ experiments
The forward model closely follows that described in Brehmer et al. [4]. We generate 5000 ensembles (drawn from the same global parameters $\theta_{\mathrm{glob}}=\{\alpha,\beta\}$), each containing 25 lensed images with varying local parameters $\theta_{\mathrm{loc}}=\{\theta_{E},\Delta x,\Delta y,q\}$. We use a ResNet-18 [12] convolutional neural network as the per-event encoder $s^{\mathrm{glob}}_{\phi}$ to extract a 128-dim embedding vector from each image; half of the features are concatenated with the true global parameters and used to condition masked autoregressive normalizing flows [20, 22] $q_{\varphi^{\prime}}^{\text{loc}}$ characterizing the local parameters for each image. The other half of the features are averaged, concatenated with the (randomly-varying at training-time) cardinality of the image dataset, and passed through a 3-layer MLP decoder, subsequently conditioning a normalizing flow $q_{\varphi}^{\text{glob}}$ modeling the global parameters of interest.

### v2

@@ Acknowledgments
Appendix A Additional details on experiments
@@ A.1 Details of ‘Simple multi-variate normal likelihood’ experiments
The estimators are trained using the AdamW (Loshchilov & Hutter, 2019; Kingma & Ba, 2014) optimizer with initial learning rate $3\times 10^{-4}$ and cosine annealing over 100 epochs. The checkpoint corresponding to the lowest validation loss is used. Evaluation is performed on 500 new test samples, evaluating the posterior estimator for sequences with cardinality varied from 1 to 200 for each test sequence.
@@ A.2 Details of ‘Mixture models in particle physics: frequentist treatment’ experiments
We use the implementation of the NUTS algorithm available in PyMC5 Salvatier et al. (2016). For each inference run, two Markov Chains of 50k samples each are sampled. The first 12k samples in each chain are discarded and not used for the estimation of the posterior.
@@ A.4 Details of ‘Astrophysics example: Strong gravitational lensing’ experiments
We generate 5000 ensembles (drawn from the same global parameters $\theta_{\mathrm{glob}}=\{\alpha,\beta\}$), each containing 25 lensed images with varying local parameters $\theta_{\mathrm{loc}}=\{\theta_{E},\Delta x,\Delta y,q\}$. We use a ResNet-18 He et al. (2016) convolutional neural network as the per-event encoder $s^{\mathrm{glob}}_{\phi}$ to extract a 128-dim embedding vector from each image; half of the features are concatenated with the true global parameters and used to condition masked autoregressive normalizing flows Papamakarios et al. (2017); Rezende & Mohamed (2015) $q_{\varphi^{\prime}}^{\text{loc}}$ characterizing the local parameters for each image. The other half of the features are averaged, concatenated with the (randomly-varying at training-time) cardinality of the image dataset, and passed through a 3-layer MLP decoder, subsequently conditioning a normalizing flow $q_{\varphi}^{\text{glob}}$ modeling the global parameters of interest.
@@
Appendix B Motivation for loss function for joint posterior inference
We motivate the loss function used in equation 7, which factorizes the joint posterior over the local and global parameters $z_{i}$ and $\theta$ into a global posterior given the set $\{x\}$ and a product of posteriors over local parameters additionally conditioned on the global parameters:
Given a hierarchical model with global parameters $\theta$ and local parameters $z_{i}$, and associated observed data $\{x\}$, we aim to demonstrate the factorization
$p(\theta,\{z_{i}\}|\{x\})\propto p(\theta|\{x\})\prod_{i}p(z_{i}|\theta,x_{i}).$ (16)
We make two key assumptions, given the nature of common hierarchical models. (a) Given the observations $\{x\}$, the global parameters $\theta$ are conditionally independent of the local parameters $\{z_{i}\}$, which gives us
$p(\theta,\{z_{i}\}|\{x\})=p(\theta|\{x\})\,p(\{z_{i}\}|\theta,\{x\}),$ (17)
and (b) Each local parameter $z_{i}$ is conditionally independent of any other $z_{j}$ given $\theta$ and its respective observation $x_{i}$, which gives
$p(\{z_{i}\}|\theta,\{x\})=\prod_{i}p(z_{i}|\theta,x_{i}).$ (18)
Substituting the result from (b) into the expression from (a), we obtain our targeted factorized form:
$p(\theta,\{z_{i}\}|\{x\})=p(\theta|\{x\})\prod_{i}p(z_{i}|\theta,x_{i})$ (19)
justifying the loss function used in the hierarchical model.